# Jl. Ganesha 10, Bandung 40132, Indonesia

### v1

@@ Abstract
Galaxy clusters are the largest virialized structures in the Universe and are predominantly dominated by dark matter. The hydrostatic mass and the mass obtained from gravitational lensing measurements generally differ, a discrepancy known as the hydrostatic mass bias. In this work, we derive the hydrostatic mass of galaxy clusters within the framework of Rastall gravity and investigate its implications under two scenarios: (i) the absence of dark matter and (ii) the existence of dark matter. In the first scenario, Rastall gravity effectively reduces the hydrostatic mass, bringing it closer to the observed baryonic mass. The best linear fit yields a slope $\mathbf{M}=1.07\pm 0.11$, indicating a near one-to-one correspondence between the two masses. In the second scenario, Rastall gravity helps to alleviate the hydrostatic mass bias. The linear fit between the Rastall hydrostatic mass and the observed lensing mass results in a best-fit slope $\mathbf{M}=1.01\pm 0.16$, which is very close to unity. These results suggest that Rastall gravity provides a statistically favorable framework for addressing mass discrepancies in galaxy clusters.
@@ 1 Introduction
The mass of galaxy clusters can be determined using three main methods: velocity dispersion measurements, X-ray or Sunyaev-Zel’dovich (SZ) observations under the assumption of hydrostatic equilibrium, and gravitational lensing. Except for gravitational lensing, these methods estimate the cluster mass within the framework of the Newtonian approximation [4, 6]. The reviews of the aforementioned methods are available in Refs. [2, 7, 8, 9, 10, 11, 12, 13, 14, 15].
It should be noted that the lensing mass of galaxy clusters is systematically higher than the hydrostatic mass. This discrepancy is commonly referred to as the hydrostatic mass bias [16]. In terms of mass composition, galaxy clusters are dominated by dark matter (DM) [17], which can account for approximately $80-90\%$ of their total mass [18]. Based on computational simulations, galaxy clusters are identified as the largest gravitationally bound DM halos in the large-scale structure of the Universe [19]. These facts highlight the challenges in accurately determining the masses of galaxy clusters.
@@
Another attempt to utilize modified theories of gravity to calculate the masses of galaxy clusters was carried out by the authors in Ref. [4]. They employed Eddington-inspired Born–Infeld (EiBI) theory, beyond Horndeski gravity (BHG), and Newtonian gravity modified by the generalized uncertainty principle (GUP). Among these theories, EiBI theory yields the best linear fit with $\mathbf{M}=0.126\pm 0.086$. This value is significantly lower than that reported by Rahvar and Mashhoon and is also much farther from unity. Consequently, EiBI theory, BHG, and Newtonian gravity modified by the GUP cannot satisfactorily account for the masses of galaxy clusters without involving DM.
@@
The aims of this work are twofold. First, we construct the formalism for the hydrostatic mass of galaxy clusters within the framework of Rastall gravity and determine the best-fit value of $\mathbf{M}$, following the methodology adopted in Refs. [4, 22, 23], with the goal of assessing whether a viable mass modeling can be achieved without invoking dark matter. Second, we determine the best-fit value of $\mathbf{M}$ in order to investigate whether the hydrostatic mass bias can be alleviated within this framework, an issue that was not resolved in Ref. [16]. We find that the Rastall parameter has a nontrivial influence on the hydrostatic mass of galaxy clusters, and that the resulting best-fit values of $\mathbf{M}$ are close to unity, with $\mathbf{M}=1$ lying within the associated uncertainty range.
The structure of this paper is as follows. Section 2 presents the mathematical formalisms employed in this work, including Rastall gravity, the standard hydrostatic formulation for determining the mass of galaxy clusters, and the hydrostatic mass formulation within the framework of Rastall gravity. Section 3 presents the numerical results, followed by a discussion of the corresponding implications. Finally, our conclusions are summarized in Section 4.
@@ 2.1 Rastall gravity
where $\lambda$ denotes the Rastall free parameter and $\mathcal{R}$ is the Ricci scalar. It is evident that in the limit $\lambda=0$, Rastall gravity is reduced to GR. Thus, $\lambda$ presents the deviation of Rastall gravity from GR.
@@
Here, $G^{\mu}_{\nu}=R^{\mu}_{\nu}-\frac{1}{2}\mathcal{R}\delta^{\mu}_{\nu}$ denotes the Einstein tensor, $\mathcal{R}^{\mu}_{\nu}$ denotes the Ricci tensor, $\delta^{\mu}_{\nu}$ is the Kronecker delta, and $\kappa$ is the constant of the field equation. For 4-dimensional spacetime, Eq. (3) gives
@@
where $T$ denotes the trace of EMT. Consequently, the Rastall field equation now becomes
@@
where $\mathcal{T}^{\mu}_{\nu}$ is the effective modified EMT that satisfies
$\mathcal{T}^{\mu}_{\nu}=T^{\mu}_{\nu}-\delta^{\mu}_{\nu}\left(\frac{\kappa\lambda}{4\kappa\lambda-1}T\right).$ (6)
@@
From Eqs. (4), (7), and (8), we have an insight that $\kappa\lambda=1/4$, $\gamma=-1/2$, and $\kappa\gamma=1/4$ are forbidden.
@@ 2.2 Standard hydrostatic mass calculation
where $p$ is the pressure, $\rho$ denotes the density of the distribution, $m(r)$ is the mass of the clusters enclosed within a radius $r$. Here, the equation of state (EOS) of the galaxy clusters is assumed to be satisfied by the EOS of the ideal gas. Such EOS is employed by Ref. [4, 6]. For the ideal gases, we have [4]
@@
$m(r)$ $=$ $-\frac{k_{B}r^{2}}{m_{p}\mu G\rho}\left(\tau\frac{d\rho}{dr}+\rho\frac{d\tau}{dr}\right)$
$m(r)$ $=$ $-\frac{k_{B}r\tau}{m_{p}\mu G}\left(\frac{r}{\rho}\frac{d\rho}{dr}+\frac{r}{\tau}\frac{d\tau}{dr}\right).$ (11)
It is widely known that
$\left(\frac{r}{\rho}\frac{d\rho}{dr}+\frac{r}{\tau}\frac{d\tau}{dr}\right)=\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln\tau}{d\ln r}\right).$ (12)
By substituting Eq. (12) into Eq. (11), one can obtain
@@
Following Ref. [22], the hot gas density of galaxy clusters is given by
@@
$n_{e}(r)n_{p}(r)=\frac{(r/r_{c})^{-\alpha^{\prime}}}{(1+r^{2}/r_{r}^{2})^{3\beta-\alpha^{\prime}/2}}\frac{n_{0}^{2}}{(1+r^{\gamma}/r_{s}^{\gamma})^{\varepsilon/\gamma}}+\frac{n_{0}^{\prime 2}}{(1+r^{2}/r_{c}^{\prime 2})^{2\beta^{\prime}}}.$ (16)
Table 1 represents the parameter related to the density for ten galaxy clusters.
@@
For the outside region of the central coolong zone,
@@
The entity $\frac{d\ln\rho}{d\ln r}$ in Eq. (14) is given by
$\frac{d\ln\rho(r)}{d\ln r}={}$ $\frac{1}{H(r)}\Bigg[-n_{0}^{2}\alpha^{\prime}\left(\frac{r}{r_{c}}\right)^{-\alpha^{\prime}}$ (21)
$\times\left(1+\frac{r^{2}}{r_{c}^{2}}\right)^{0.5\alpha^{\prime}-3\beta^{\prime}}\left(1+r^{\gamma}r_{s}^{-\gamma}\right)^{-\frac{\varepsilon}{\gamma}}$
$+n_{0}^{2}(\alpha^{\prime}-6\beta)\left(\frac{r}{r_{c}}\right)^{-\alpha^{\prime}+2}$
$\times\left(1+\frac{r^{2}}{r_{c}^{2}}\right)^{-1+0.5\alpha^{\prime}-3\beta}\left(1+r^{\gamma}r_{s}^{-\gamma}\right)^{-\frac{\varepsilon}{\gamma}}$
$-6n_{0}^{\prime 2}\beta^{\prime}\left(\frac{r}{r_{c}^{\prime}}\right)^{2}\left(1+\frac{r^{2}}{r_{c}^{\prime 2}}\right)^{-1-3\beta^{\prime}}$
$-n_{0}^{2}\varepsilon\left(\frac{r}{r_{c}}\right)^{-\alpha^{\prime}}\left(\frac{r}{r_{s}}\right)^{\gamma}$
$\times\left(1+\frac{r^{2}}{r_{c}^{2}}\right)^{0.5\alpha^{\prime}-3\beta}\left(1+r^{\gamma}r_{s}^{-\gamma}\right)^{-1-\frac{\varepsilon}{\gamma}}\Bigg].$
where
$H(r)={}$ $2\Bigg[n_{0}^{\prime 2}\left(1+\frac{r^{2}}{r_{c}^{\prime 2}}\right)^{-3\beta^{\prime}}$ (22)
$\quad+n_{0}^{2}\left(\frac{r}{r_{c}}\right)^{-\alpha^{\prime}}\left(1+\frac{r^{2}}{r_{c}^{2}}\right)^{0.5\alpha^{\prime}-3\beta}$
$\quad\times\left(1+r^{\gamma}r_{s}^{-\gamma}\right)^{-\frac{\varepsilon}{\gamma}}\Bigg].$
@@ 2.3 Hydrostatic mass of galaxy clusters in Rastall gravity
$ds^{2}=-e^{2\nu(r)}dt^{2}+e^{2\lambda(r)}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta,d\phi^{2},$ (23)
@@
$T^{\mu}_{\ \nu}=\begin{pmatrix}-\rho&0&0&0\\ 0&p&0&0\\ 0&0&p&0\\ 0&0&0&p\end{pmatrix}$. (24)
@@
Recall that, in GR, the hydrostatic equilibrium equation in the Newtonian regime can be recovered from the TOV equation by assuming $\rho\gg p$ and $r\gg m$. By adopting the same procedure and also considering $\frac{d\rho}{dp}=\frac{\rho}{r}\frac{d\ln\rho}{d\ln r}\left(\frac{dp}{dr}\right)^{-1}$, we obtain the Newtonian-limit expression of Eq. (25), i.e.
@@
$\frac{dp}{dr}=\left[-\frac{\rho}{r^{2}}\left\{m+\frac{\kappa\lambda}{4\kappa\lambda-1}\left(\frac{\kappa r^{3}\rho}{2}+r\frac{d\ln\rho}{d\ln r}\right)\right\}\right]\frac{4\kappa\lambda-1}{\kappa\lambda-1}.$ (27)
@@
$m(r)=-\frac{k_{B}}{\mu m_{p}}\frac{\kappa\lambda-1}{4\kappa\lambda-1}\tau(r)r\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln\tau}{d\ln r}\right)-\frac{\kappa\lambda}{4\kappa\lambda-1}\left(\frac{\kappa r^{3}\rho}{2}+\frac{c^{2}r}{G}\frac{d\ln\rho}{d\ln r}\right).$ (28)
Eq. (28) is the Rastall version of the hydrostatic mass of galaxy clusters. When $\lambda=0$, the equation is reduced to Eq. (13).
(a) (a) Linear fit to the relation between the hydrostatic masses according to standard non-Rastall gravity and the observed baryonic masses, (b) the likelihood function for parameter $\mathbf{M}$ in standard non-Rastall gravity, (c) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=9.55\times 10^{-2}$ and the observed baryonic masses, (d) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=9.55\times 10^{-2}$, (e) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=1.14\times 10^{-1}$ and the observed baryonic masses, and (f) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=1.14\times 10^{-1}$.
(b) (a) Linear fit to the relation between the hydrostatic masses according to standard non-Rastall gravity and the observed lensing masses, (b) the likelihood function for parameter $\mathbf{M}$ in standard non-Rastall gravity, (c) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=1.49\times 10^{-3}$ and the observed baryonic masses, (d) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=1.49\times 10^{-3}$, (e) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=1.69\times 10^{-3}$ and the observed baryonic masses, and (f) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=1.69\times 10^{-3}$.
3 Numerical results and discussion

### v2

# Impact of Rastall gravity on hydrostatic mass of galaxy clusters

@@ Abstract
Galaxy clusters are the largest virialized structures in the Universe and are predominantly dominated by dark matter. The hydrostatic mass and the mass obtained from gravitational lensing measurements generally differ, a discrepancy known as the hydrostatic mass bias. In this work, we derive the hydrostatic mass of galaxy clusters within the framework of Rastall gravity. We consider two scenarios: (i) the absence of dark matter and (ii) the presence of dark matter. In both cases, we constrain the Rastall parameter in the cluster-scale using observational data. In the first scenario, Rastall gravity effectively reduces the hydrostatic mass, bringing it closer to the observed baryonic mass. The best linear fit yields a slope $\mathbf{M}=1.07\pm 0.11$, indicating a near one-to-one correspondence between the two masses. In the second scenario, Rastall gravity helps to alleviate the hydrostatic mass bias. The linear fit between the Rastall hydrostatic mass and the observed lensing mass results in a best-fit slope $\mathbf{M}=0.99\pm 0.26$, which is very close to unity. We also calculate the goodness-of-fit for every fit. The statistical evaluations indicate that Rastall gravity provides a viable phenomenological framework that can improve certain aspects of the mass discrepancy problem at the level of scaling relations. However, it does not universally outperform other modified gravity model, when evaluated using standard goodness-of-fit criteria.
keywords:
Galaxy cluster , Hydrostatic mass , Rastall gravity
journal: Physics Letters B
[…]
@@ 1 Introduction
The mass of galaxy clusters can be determined using three main methods: velocity dispersion measurements, X-ray or Sunyaev-Zel’dovich (SZ) observations under the assumption of hydrostatic equilibrium, and gravitational lensing. Except for gravitational lensing, these methods estimate the cluster mass within the framework of the Newtonian approximation [4, 6]. The reviews of the aforementioned methods are available in Refs. [2, 7, 8, 9, 10, 11, 12, 13, 14, 15]. The hydrostatic mass approach has been commonly used in the studies of galaxy clusters [16, 17, 18, 19, 20, 21]. Note that there is another method to calculate mass of galaxy clusters, i.e. richness proxy. Here, richness is defined as the number of bright cluster galaxies [22]. The details of the richness proxy can be found in Refs. [22, 23, 24, 25, 26]
It should be noted that generally, the lensing mass of galaxy clusters is systematically higher than the hydrostatic mass. This discrepancy is commonly referred to as the hydrostatic mass bias [27]. In terms of mass composition, galaxy clusters are dominated by dark matter (DM) [28], which can account for approximately $80-90\%$ of their total mass [29]. Based on computational simulations, galaxy clusters are identified as the largest gravitationally bound DM halos in the large-scale structure of the Universe [30].
@@
Another attempt to utilize modified theories of gravity to calculate the masses of galaxy clusters was carried out by the authors in Ref. [4]. They employed Eddington-inspired Born–Infeld (EiBI) theory, beyond Horndeski gravity (BHG), and Newtonian gravity modified by the generalized uncertainty principle (GUP). The use of GUP in the calculation of the mass of galaxy clusters can also be found in Ref. [35]. Among these theories, EiBI theory yields the best linear fit with $\mathbf{M}=0.126\pm 0.086$. This value is significantly lower than that reported by Rahvar and Mashhoon and is also much farther from unity. Consequently, EiBI theory, BHG, and Newtonian gravity modified by the GUP cannot satisfactorily account for the masses of galaxy clusters without involving DM.
@@
The aims of this work are twofold. First, we construct the formalism for the hydrostatic mass of galaxy clusters within the framework of Rastall gravity. Second, we determine the best-fit values of in two scenarios: (i) the absence of DM, and (ii) the presence of DM; in order to constrain the Rastall parameter at cluster scales in both cases. In the first scenario, the Rastall-modified hydrostatic mass is compared with the observed baryonic mass of galaxy clusters. In the second scenario, the Rastall-modified hydrostatic mass is compared with the lensing mass data to examine whether Rastall gravity can alleviate the hydrostatic mass bias. We note that alleviating the discrepancy between hydrostatic and lensing mass estimates remains an open issue, as it was not resolved in Ref. [27]. We find that the Rastall parameter has a nontrivial influence on the hydrostatic mass of galaxy clusters, and that the resulting best-fit values of $\mathbf{M}$ are close to unity, with $\mathbf{M}=1$ lying within the associated uncertainty range.
The structure of this paper is as follows. Section 2 presents the mathematical formalisms employed in this work, including Rastall gravity, the standard hydrostatic formulation for determining the mass of galaxy clusters, and the hydrostatic mass formulation within the framework of Rastall gravity. Section 3 presents the numerical results and statistical evaluations on the results, followed by a discussion of the corresponding implications. Finally, our conclusions are summarized in Section 4.
@@ 2.1 Rastall gravity
where $\lambda$ denotes the Rastall free parameter and $\mathcal{R}$ is the Ricci scalar. As a consequence of Eq. (2), the EMT is no longer conservative. It is evident that in the limit $\lambda=0$, Rastall gravity is reduced to GR. Thus, $\lambda$ represents the deviation of Rastall gravity from GR.
@@
Here, $G^{\mu}_{\nu}=\mathcal{R}^{\mu}_{\nu}-\frac{1}{2}\mathcal{R}\delta^{\mu}_{\nu}$ denotes the Einstein tensor, $\mathcal{R}^{\mu}_{\nu}$ denotes the Ricci tensor, $\delta^{\mu}_{\nu}$ is the Kronecker delta, and $\kappa$ is the constant of the field equation. For 4-dimensional spacetime, Eq. (3) gives
@@
where $T$ denotes the trace of EMT. Consequently, the Rastall non-conservation law can be written as
${\color[rgb]{0,0,0}\nabla_{\mu}T^{\mu\nu}=\frac{\kappa\lambda}{4\kappa\lambda-1}\nabla^{\nu}T}.$ (5)
Furthermore, Rastall field equation now becomes
@@
where $\mathcal{T}^{\mu}_{\nu}$ is the effective modified EMT that satisfies the definition
$\mathcal{T}^{\mu}_{\nu}{\color[rgb]{0,0,0}\equiv}T^{\mu}_{\nu}-\delta^{\mu}_{\nu}\left(\frac{\kappa\lambda}{4\kappa\lambda-1}T\right).$ (7)
@@
From Eqs. (4), (8), and (9), we have an insight that $\kappa\lambda=1/4$, $\gamma=-1/2$, and $\kappa\gamma=1/4$ are forbidden, since the denominator will be zero. In this work, the values of $\gamma$, $\kappa$, and $\lambda$ are shown in Table 1. It is worth noting that those values are the parameters that result in the best-fits.
Table 1: Values of $\gamma$, $\kappa$, and $\lambda$ used in this work.
$\gamma$ $\kappa$ $\lambda$
$4.00$ $0.89\pi$ $9.55\times 10^{-2}$
$4.95$ $0.73\pi$ $1.14\times 10^{-1}$
$3.50\times 10^{-2}$ $8.60\pi$ $1.14\times 10^{-3}$
$4.10\times 10^{-2}$ $8.71\pi$ $1.29\times 10^{-3}$
Note that the original Rastall’s theory [36] was derived by re-questioning the conservation of the EMT, rather than being derived from the action. Despite this fact, Ref. [68] successfully derived the Rastall-type field equation from $f(\mathcal{R},T)$ gravity. In $f(\mathcal{R},T)$ gravity, the action writes [69]
${\color[rgb]{0,0,0}S=\frac{1}{2\kappa}\int d^{4}x\sqrt{-g}\bigg(f(\mathcal{R},T)+\mathscr{L}_{m}\bigg).}$ (10)
Here, $g$ denotes the determinant of the metric tensor, and $\mathscr{L}_{m}$ is the Lagrangian of matter.
In Ref. [68], a specific function $f(\mathcal{R},T)$ is chosen, i.e.
${\color[rgb]{0,0,0}f(\mathcal{R},T)=\mathcal{R}+\kappa\varsigma T}.$ (11)
Here, $\varsigma$ is the free parameter of the $f(\mathcal{R},T)$ gravity theory. By deriving the action containing the function shown in Eq. (11), one obtains the Rastall-type field equation, i.e.
${\color[rgb]{0,0,0}G_{\mu\nu}=\frac{\kappa}{1-\varsigma}\bigg[(1-\varsigma)T_{\mu\nu}+\varsigma\bigg(\frac{g_{\mu\nu}}{2}T-g^{\alpha\beta}\frac{\delta T_{\alpha\beta}}{\delta g^{\mu\nu}}\bigg)\bigg].}$ (12)
For the static spherically symmetric perfect fluid EMT $T_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}+pg_{\mu\nu}$, where $\rho$, $p$, and $u_{\mu}$ denote energy density distribution of matter, pressure, and 4-velocity, respectively; we can easily find that
${\color[rgb]{0,0,0}g^{\alpha\beta}\frac{\delta T_{\alpha\beta}}{\delta g^{\mu\nu}}=-2(\rho+p)u_{\mu}u_{\nu}-pg_{\mu\nu}.}$ (13)
By defining $\mathcal{T}^{*}=g^{\mu\nu}\bigg(g_{\mu\nu}T-2g^{\alpha\beta}\frac{\delta T_{\alpha\beta}}{\delta g^{\mu\nu}}\bigg)$, one obtains
${\color[rgb]{0,0,0}\nabla_{\mu}T^{\mu\nu}=\frac{\varsigma}{2(\varsigma-1)}\nabla^{\nu}\mathcal{T}^{*},}$ (14)
which is similar to Eq. (5) and exhibits the same interpretation, i.e., the non-conservation of EMT.
Within the weak-field regime, using Eq. (13) and assuming a static, spherically symmetric perfect fluid with $\rho\gg p$ and $r\gg m$, we obtain $\mathcal{T}^{*}=-5\rho$. By choosing
${\color[rgb]{0,0,0}\varsigma=-\frac{2}{5\bigg[1-\frac{2\kappa\lambda}{5(4\kappa\lambda-1)}\bigg]}\frac{\kappa\lambda}{4\kappa\lambda-1},}$ (15)
and recalling that $T=-\rho$ in the weak-field regime, we obtain
${\color[rgb]{0,0,0}\nabla_{\mu}T^{\mu\nu}}{\color[rgb]{0,0,0}=}$ ${\color[rgb]{0,0,0}-\frac{\kappa\lambda}{4\kappa\lambda-1}\nabla^{\nu}\rho,}$
${\color[rgb]{0,0,0}=}$ ${\color[rgb]{0,0,0}\frac{\kappa\lambda}{4\kappa\lambda-1}\nabla^{\nu}T,}$ (16)
which is nothing but the Rastall non-conservation law shown by Eq. (5). At this stage, we have shown that, for a static spherically symmetric perfect fluid, Rastall gravity in the weak-field limit can be effectively derived from $f(\mathcal{R},T)$ gravity.
@@ 2.2 Standard hydrostatic mass calculation
where $m(r)$ is the mass of the clusters enclosed within a radius $r$. Here, the equation of state (EOS) of the galaxy clusters is assumed to be satisfied by the EOS of the ideal gas. Such EOS is employed by Ref. [4, 6]. For the ideal gases, we have [4]
@@
Now we proceed to the profiles related to the galaxy clusters. According to Ref. [33], the temperature of the intracluster gas is typically of the order of keV. At such temperatures, the hot plasma predominantly emits X-rays through thermal bremsstrahlung (free–free emission). In addition, line emission from ionized heavy elements also contributes to the overall radiation. The total emitted radiation is approximately proportional to the product of the electron number density $n_{e}$ and proton number density $n_{p}$. Following Ref. [33], the hot gas density of galaxy clusters is given by
@@
$n_{e}(r)n_{p}(r)$ $=$ $\frac{(r/r_{c})^{-\alpha^{\prime}}}{(1+r^{2}/r_{c}^{2})^{3\beta-\alpha^{\prime}/2}}\frac{n_{0}^{2}}{(1+r^{\gamma}/r_{s}^{\gamma})^{\varepsilon/\gamma}}$ (22)
$+\frac{n_{0}^{\prime 2}}{(1+r^{2}/r_{c}^{\prime 2})^{2\beta^{\prime}}}.$
Table 2 represents all parameters needed in Eq. (22).
@@
In this sequence of calculations, each cluster is considered to be a spherical configuration of matter with an effective radius of $r_{500}$. Here, $r_{500}$ is defined to be the cluster radius within which the average overdensity is 500 times the critical density of the universe at the redshift of the cluster in the DM model [33].
@@
For the outside region of the central cooling zone,
@@
In order to calculate the hydrostatic mass by using Eq. (20), we need the radial derivatives of the three-dimensional gas density and temperature profiles. The entity $\frac{d\ln\rho}{d\ln r}$ in Eq. (20) is given by
More details on the standard calculation of hydrostatic mass can be referred to Refs. [33, 71]. The values of the parameters needed in the calculation of $\frac{d\ln\rho}{d\ln r}$ are presented in Table 2, while the ones needed in the calculation of $\frac{d\ln\tau}{d\ln r}$ are presented in Table 3.
@@ 2.3 Hydrostatic mass of galaxy clusters in Rastall gravity
$ds^{2}=-e^{2\nu(r)}dt^{2}+e^{2\lambda(r)}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2},$ (30)
@@
$T^{\mu}_{\ \nu}=\begin{pmatrix}-\rho{\quad}&0{\quad}&0{\quad}&0\\[3.44444pt] 0{\quad}&p{\quad}&0{\quad}&0\\[3.44444pt] 0{\quad}&0{\quad}&p{\quad}&0\\[3.44444pt] 0{\quad}&0{\quad}&0{\quad}&p\end{pmatrix}$. (31)
@@
Recall that, in GR, the hydrostatic equilibrium equation in the Newtonian regime can be recovered from the TOV equation by assuming $\rho\gg p$ and $r\gg m$, as has been mentioned before. By adopting the same procedure and also considering $\frac{d\rho}{dp}=\frac{\rho}{r}\frac{d\ln\rho}{d\ln r}\left(\frac{dp}{dr}\right)^{-1}$, we obtain the Newtonian-limit expression of Eq. (32), i.e.
@@
$\frac{dp}{dr}$ $=$ $\left[-\frac{\rho}{r^{2}}\left\{m+\frac{\kappa\lambda}{4\kappa\lambda-1}\left(\frac{\kappa r^{3}\rho}{2}+r\frac{d\ln\rho}{d\ln r}\right)\right\}\right]$ (34)
$\times\frac{4\kappa\lambda-1}{\kappa\lambda-1}.$
@@
$m(r)$ $=$ $-\frac{k_{B}}{\mu m_{p}}\frac{\kappa\lambda-1}{4\kappa\lambda-1}\tau(r)r\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln\tau}{d\ln r}\right)$ (35)
$-\frac{\kappa\lambda}{4\kappa\lambda-1}\left(\frac{\kappa r^{3}\rho}{2}+\frac{c^{2}r}{G}\frac{d\ln\rho}{d\ln r}\right).$
Eq. (35) is the Rastall version of the hydrostatic mass of galaxy clusters. The entities $\tau$, $\frac{d\ln\rho}{d\ln r}$, and $\frac{d\ln\tau}{d\ln r}$ can be calculated by using Eqs. (26), (LABEL:dlnrho), and (LABEL:dlntau), respectively. When $\lambda=0$, the equation is reduced to Eq. (19).
Figure 1: (a) Linear fit to the relation between the hydrostatic masses according to standard non-Rastall gravity and the observed baryonic masses, (b) the likelihood function for parameter $\mathbf{M}$ in standard non-Rastall gravity, (c) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=9.55\times 10^{-2}$ and the observed baryonic masses, (d) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=9.55\times 10^{-2}$, (e) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=1.14\times 10^{-1}$ and the observed baryonic masses, and (f) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=1.14\times 10^{-1}$.
Figure 2: (a) Linear fit to the relation between the hydrostatic masses according to standard non-Rastall gravity and the observed lensing masses, (b) the likelihood function for parameter $\mathbf{M}$ in standard non-Rastall gravity, (c) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=1.14\times 10^{-3}$ and the observed baryonic masses, (d) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=1.14\times 10^{-3}$, (e) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=1.29\times 10^{-3}$ and the observed baryonic masses, and (f) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=1.29\times 10^{-3}$.
3 Numerical results, statistical evaluations, and discussion